OpenStudy (anonymous):

Given: f(x)= 7x^3 + 2 A. identify the function as even, odd, or neither. B. Describe the end behavior of the function.

OpenStudy (freckles):

Find f(-x)

OpenStudy (freckles):

\[f(x)=7x^3+2 \\ f(-x)=7(-x)^3+2\] what is the cube of (-1)?

OpenStudy (anonymous):

1?

OpenStudy (freckles):

\[(-x)^3=(-1x)^3=(-1)^3(x)^3=(-1)^3x^3=(-1 \cdot -1 \cdot -1)x^3\]

OpenStudy (freckles):

-1 times -1 is 1

OpenStudy (freckles):

but -1 times -1 times -1 is -1

OpenStudy (anonymous):

ops i squared it on accident

OpenStudy (freckles):

\[(-x)^3=-1x^3 \text{ or just } -x^3\]

OpenStudy (freckles):

but anyways this means \[f(-x)=-7x^3+2\]

OpenStudy (freckles):

But this isn't the same as 7x^3+2 or the opposite right?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

if it was the same as the function we started with would be even if it was the opposite of the function we started with it would be odd so yes you are right it is neither

OpenStudy (freckles):

There is a trick to doing polynomials quicker if you want to know

OpenStudy (anonymous):

sure

OpenStudy (freckles):

If all the powers in the polynomial are odd, then the function is odd. If all the powers in the polynomial are even, then the function is even. If you have a mixture of evens and odds powers, then the function is neither odd or even. Now a lot of people get confused when they don't see a variable next to one of the numbers. But you can think of your function as f(x)=7x^3+2x^0

OpenStudy (freckles):

3 is odd 0 is even so f is neither since you have a mixture of even and odd powers

OpenStudy (freckles):

What I just said works for any polynomial... f(x)=3 is even because again f(x)=3x^0 and 0 is even f(x)=x is odd because f(x)=x^1 and 1 is odd f(x)=x^2+3 is even because f(x)=x^2+3x^0 and 2,0 are even

OpenStudy (anonymous):

interesting

OpenStudy (freckles):

Do you know what I mean by a polynomial?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[P(x)=a_0x^0+a_1x^1+a_2x^2+a_3x^3+a_4x^4+a_5x^5+ \cdots a_nx^{n}\] where n is integer

OpenStudy (freckles):

\[P_1(x)=a_0x^0+a_2x^2+a_4x^4+a_6x^6 \text{ is even } \\ P_2(x)=a_1x^1+a_3x^3+a_5x^5 \text{ is odd }\]

OpenStudy (freckles):

The function I defined as P_1 is even because 0,2,4,6 are even The function I defined as P_2 is odd because 1,3,5 are odd

OpenStudy (anonymous):

I think i got it

OpenStudy (freckles):

lol ok

OpenStudy (freckles):

So knowing that it is odd that should tell you about the opposite ends of our function.

OpenStudy (freckles):

Oops sorry I was thinking about a different function.

OpenStudy (freckles):

We are looking at f(x)=7x^3+2

OpenStudy (freckles):

and that function had neither symmetry

OpenStudy (anonymous):

so what does that mean for the end?

OpenStudy (freckles):

What are you allowed to do? I can do limits with you or you could graph in your calculator.

OpenStudy (aum):

Are you familiar with the graph of y = x^3 ?

OpenStudy (freckles):

The ends means like what is happening as we go view the graph for really large values of x or really small values of x

OpenStudy (anonymous):

oh okay wouldnt it be kind of small values?

OpenStudy (aum):